OpenStudy (anonymous):

Write the equation of a line in point-slope form that has a slope of -1 and passes through the point (-2, 5)

OpenStudy (the_fizicx99):

Y-y1=m(x-x1) y-5=-1(x-(-2)) Solve from there?

OpenStudy (anonymous):

can you explain it more?

OpenStudy (the_fizicx99):

The formula is y-y1=m(x-1) Plug in those values into the formula M or "slope" = -1 (-2,5) --> (x,y)

OpenStudy (anonymous):

got it! thanks!

OpenStudy (the_fizicx99):

Yw